The area of a triangle is 13.5 square meters. Find the base and height of the rectangle if its height is 6 meters greater than its base. Use an equation and the formula area of a triangle 0.5(base)(height).
Answer by checkley77(12844)

AREA=.5(B*H)
13.5=.5(B)
13.5=.5(B^2+6B)
B^2+6B=13.5/.5
B^2+6B=27
B^2+6B-27=0
(B+9)(B-3)=0
B-3=0
B=3 FOR THE BASE.
3+6=9 FOR THE HEIGHT.
PROOF:
13.5=.5(3*9)
13.5=.5*27
13.5=13.5